OpenStudy (lgg23):

With arbitrary launch angle, the range will be diluted by cos(theta): x=((v sub-o)cos(theta))t, while the vertical distance fallen is y, given by -y=((v sub-o)sin(theta))t-(1/2)g(t^2) Can you combine these equations and seek the angle theta that theoretically maximizes the range x, when launched (as in our experiment) at height h above the landing field? (Hint: this procedure is relatively easy on a horizontal surface, where y=0; there the maximum range occurs when theta=45deg.)

OpenStudy (goformit100):

post the question in mathematics section

OpenStudy (anonymous):

Now, we need to find the angle at which x is maximum. Since, at the instant particle lands, vertical distance dropped is h, we'll put y=h. So, the equation becomes \[-h = v _{0}\sin (\theta)t-\frac{ 1 }{ 2 }g t^{2}\] On eliminating t from the 1st and 3rd equation, we get \[-h \cos ^{2}\theta = \frac{ x }{ 2 }\sin 2\theta - \frac{ 1 }{ 2 }g \frac{ x^{2} }{ v _{0} ^{2}} \] Taking the derivative w.r.t theta on both sides, we'll next put \[\frac{ dx }{ d \theta } = 0\] which is the maxima condition for x. I believe you can do that.